A 50n gun fires a 1.0 gram bullet at a speed pf 2200 m/s. what is the speed pf gun recoil?
1
Expert's answer
2021-10-07T08:50:48-0400

Let's first find the mass of the gun:

"m_g = \frac{50N}{g} = \frac{50N}{10m/s^2} = 5kg"

where "g=10m\/s^2" is the gravitatinal acceleration.

The momentum of the bullet is:

"p_b = m_bv_b = 0.001kg \times 2200m/s = 2.2kg⋅m/s"

Then, according to the momentum conservation law, the momentum of the gun has the same magnitude:

"p_g = p_b = 2.2 kg\u22c5m\/s"

The speed of gun then:

"v_g = \frac{p_g}{m_g} = \frac{2.2\space kg\cdot m/s }{5kg}=0.44m/s"


Ans. 0.44m/s